You wish to test the claim that μ>12 at a level of significance of α=0.05 and are given the sample statistics mean = 12.3, deviation = 1.2, and sample size = 50. Compute the value of the standardized test statistic. A. 3.11 B. 0.98 C. 1.77 D. 2.31
Answer #1

solution Here Ho=M=12, VS Hi M>12 Given n = 12-3, s=12, n=50 Test statistic it = x - l = 4 inte - 7-4 - 12-3-12 = 1.768 1-2)5